A 0.784-g sample of an alloy containing Al and Mg is dissolved and treated to prevent interferences by the alloy’s other constituents. Aluminum and magnesium are precipitated, providing a mixed precipitate of Al(C9H6NO)3 and Mg(C9H6NO)2 that weighs 8.215 g. Igniting the precipitate converts it to a mixture of Al2O3 and MgO that weighs 0.932 g. Calculate the %w/w Al and %w/w Mg in the alloy.


1
Expert's answer
2022-04-27T06:47:04-0400

Solution:

For Al we have:

Al → Al(C9H6NO)3 → Al2O3

According to chemical scheme above:

2 × Moles of Al2O3 = Moles of Al(C9H6NO)3

Converting from moles to grams and solving yields an equation relating the grams of Al2O3 to the grams of Al(C9H6NO)3:

2 × (Mass of Al2O3 / Molar mass of Al2O3) = (Mass of Al(C9H6NO)3 / Molar mass of Al(C9H6NO)3)

Molar mass of Al2O3 is 101.96 g/mol

Molar mass of Al(C9H6NO)3 is 459.4317 g/mol

Therefore,

2 × (Mass of Al2O3 / 101.96) = (Mass of Al(C9H6NO)3 / 459.4317)

918.8634 × Mass of Al2O3 = 101.96 × Mass of Al(C9H6NO)3

Mass of Al2O3 = 0.11096 × Mass of Al(C9H6NO)3


For Mg we have:

Mg → Mg(C9H6NO)2 → MgO

According to chemical scheme above:

Moles of MgO = Moles of Mg(C9H6NO)2

Converting from moles to grams and solving yields an equation relating the grams of MgO to the grams of Mg(C9H6NO)2:

(Mass of MgO / Molar mass of MgO) = (Mass of Mg(C9H6NO)2 / Molar mass of Mg(C9H6NO)2)

Molar mass of MgO is 40.3044 g/mol

Molar mass of Mg(C9H6NO)2 is 312.605 g/mol

Therefore,

(Mass of MgO / 40.3044) = (Mass of Mg(C9H6NO)2 / 312.605)

312.605 × Mass of MgO = 40.3044 × Mass of Mg(C9H6NO)2

Mass of MgO = 0.12893 × Mass of Mg(C9H6NO)2


Mass of Al2O3 + Mass of MgO = 0.932 g

Therefore,

0.11096 × Mass of Al(C9H6NO)3 + 0.12893 × Mass of Mg(C9H6NO)2 = 0.932 g


Mass of Al(C9H6NO)3 + Mass of Mg(C9H6NO)2 = 8.215 g

0.11096 × Mass of Al(C9H6NO)3 + 0.12893 × Mass of Mg(C9H6NO)2 = 0.932 g

Multiplying the first equation by 0.11096 and subtracting the second equation gives:

−0.01797 × Mass of Mg(C9H6NO)2 = −0.020464

Mass of Mg(C9H6NO)2 = 1.1388 g


The mass of Al(C9H6NO)3 can then be calculated using the known combined mass of the two original precipitates:

Mass of Al(C9H6NO)3 = 8.215 g − Mass of Mg(C9H6NO)2 = 8.215 g − 1.1388 g = 7.0762 g

Mass of Al(C9H6NO)3 = 7.0762 g


For Al we have:

Al → Al(C9H6NO)3 → Al2O3

According to chemical scheme above:

Moles of Al = Moles of Al(C9H6NO)3

Converting from moles to grams and solving yields an equation relating the grams of Al to the grams of Al(C9H6NO)3:

(Mass of Al / Molar mass of Al) = (Mass of Al(C9H6NO)3 / Molar mass of Al(C9H6NO)3)

Molar mass of Al is 26.9815 g/mol

Molar mass of Al(C9H6NO)3 is 459.4317 g/mol

Mass of Al(C9H6NO)3 = 7.0762 g

Therefore,

Mass of Al = (Mass of Al(C9H6NO)3 × Molar mass of Al) / (Molar mass of Al(C9H6NO)3)

Mass of Al = (7.0762 g × 26.9815 g/mol) / (459.4317 g/mol) = 0.4156 g

Mass of Al = 0.4156 g

%Al = (Mass of Al / Mass of sample) × 100% = (0.4156 g / 0.784 g) × 100% = 53.01%

%Al = 53.01%


For Mg we have:

Mg → Mg(C9H6NO)2 → MgO

According to chemical scheme above:

Moles of Mg = Moles of Mg(C9H6NO)2

Converting from moles to grams and solving yields an equation relating the grams of Mg to the grams of Mg(C9H6NO)2:

(Mass of Mg / Molar mass of Mg) = (Mass of Mg(C9H6NO)2 / Molar mass of Mg(C9H6NO)2)

Molar mass of Mg is 24.305 g/mol

Molar mass of Mg(C9H6NO)2 is 312.605 g/mol

Mass of Mg(C9H6NO)2 = 1.1388 g

Therefore,

Mass of Mg = (Mass of Mg(C9H6NO)2 × Molar mass of Mg) / (Molar mass of Mg(C9H6NO)2)

Mass of Mg = (1.1388 g × 24.305 g/mol) / (312.605 g/mol) = 0.08854 g

Mass of Mg = 0.08854 g

%Mg = (Mass of Mg / Mass of sample) × 100% = (0.08854 g / 0.784 g) × 100% = 11.29%

%Mg = 11.29%


Answer:

The %w/w Al in the alloy is 53.01%

The %w/w Mg in the alloy is 11.29%